Department G had 2,280 units 25% completed at the beginning of the period, 12,900 units were completed during the period, 1,900 units were 20% completed at the end of the period, and the following manufacturing costs were debited to the departmental work in process account during the period:

Work in process, beginning of period $25,400 Costs added during period: Direct materials (12,520 units at $9) 112,680 Direct labor 127,100 Factory overhead 38,130 All direct materials are placed in process at the beginning of production, and the first-in, first-out method of inventory costing is used. What is the total cost of 2,280 units of beginning inventory that were completed during the period? a.$49,530 b.$47,630 c.$25,400 d.$43,184
Answer #1

Units started and completed = 12900 - 2280 = 10620
Equivalent units of production
Material = 2280 x (100%-100%) + 10620 + 1900 x 100% = 12520 units
Conversion cost = 2280 x (100%-25%) + 10620 + 1900 x 20% = 12710 units

Cost per unit
Material = $112680 / 12520 = $9 per unit
Conversion cost = ($127100+38130)/12710 = $13 per unit

Cost of Beginningh inventory completed = $25400 + 2280 x 75% x $13 = $47630

Answer is b. $47630